M-「-5-7 7 -5 Find formulas for the entries of Mn, where n is a positive integer. (Your formulas should not contain complex numbers.) Mn =
Answer #1

Solved using basic knowledge of linear algebra

M = -5-7] L7 - 5 Let a be the eigenvalue FM IM-XI1=0 -5 2 - 7 =0 T7 -5-1 -> (-5-7² +40=0 x ² + 10x + 74=0 d = -10 + V 100 - 2-fi-7 R, (e) 7-ti RR 17 -ti R, +7 L 7 - 7e til Li-i -> x = ing let = 1 = x = iWhen 2 = -5-ti Letaz (ori] be its eigen vector A-d₂I) 5 = [-5-6-5-22) - 7-5 [x] - 1:3 Lote | R, (A) II ir Ti i 10 cel => x=-1Binomial expansion (a+b)n = nan-kbk - m1= -4 Galo i (-5+7in - i (-5-7iNTO Ti i LC-5+7in (-5-7in 1 21) L-1 is H i(-5+7)+i (-S-KE & S (-5) 1-2h 7 2K (w 2k (2) whese th] ecs the greatest integes function on ..Min - i nc (-577-22 2h Dhal 22nc(-s) 92-9672Consider MM 1/2 = -(-5+76)n + 65-72 = - ne (-5)n-k (70k + nG (5) B-k (-70k M.2 Inc (-5)n-k (72) K (-1+(-k) - ② K = 0 Thus Mr- Ri mn M22 J 1/2 M Mn. 1/2] in n-26 (-5) (-DR 7 2K -2k-1 K (-1) 2h+ 7 2kt ko 69127 n-2k-1 0-2h -5) 27-2560572 - K2k 2k1 K=0